# Statistically-Guided Dual-Domain Meta-Learning with Adaptive Multi-Prototype Aggregation for Distributed Fiber Optic Sensing

### v1

@@ Abstract
Distributed Fiber Optic Sensing (DFOS) has shown strong potential in perimeter security due to its capability of monitoring vibration events across long distances with fine spatial resolution. However, practical DFOS systems face three critical challenges: (1) signal patterns of the same activity vary drastically under different fiber deployment types (e.g., underground, wall-mounted), causing domain shift; (2) labeled data in new deployment scenarios is often scarce or entirely unavailable, limiting model adaptability; and (3) even within source domains, data scarcity makes it difficult to capture intra-class diversity for robust learning.
To address these challenges, we propose a novel meta-learning framework, DUPLE, for cross-deployment DFOS activity identification. First, a dual-domain multi-prototype learner fuses temporal and frequency domain features, enhancing the model’s generalization ability under signal distribution shifts. Second, a Statistical Guided Network (SGN) infers domain importance and prototype sensitivity from raw statistical features, providing data-driven prior information for learning in unlabeled or unseen domains. Third, a query-aware prototype aggregation module adaptively selects and combines relevant prototypes, thereby improving classification performance even with limited data.
Extensive experiments on cross-deployment DFOS datasets demonstrate that our method significantly outperforms baseline approaches in domain generalization settings, enabling robust event recognition across diverse fiber configurations with minimal labeled data.
@@ 1 Introduction
  • Dual-domain multi-prototype meta-learner. We learn time-domain and frequency-domain embeddings and perform per-class multi-prototype clustering in each view; class evidence is combined via a soft-OR (log-sum-exp) rule, yielding expressive, deployment-robust class representations.
  • Statistical Guidance Network (SGN). From raw statistical descriptors, SGN infers domain importance and prototype sensitivity, providing data-driven priors that coordinate the two views and stabilize learning under unlabeled or unseen deployments.
  • Query-aware adaptive aggregation. In the collaborative decision stage, each query uses its SGN guidance to form a key and applies attention over the class’s prototype sets, selecting and weighting the most relevant prototypes before producing the final logits.
Together, these designs align the training objective with cross-deployment use, inject sample-wise statistical priors to guide dual-domain fusion, and enable instance-level adaptation at decision time—delivering stronger generalization with minimal data requirements.
@@ 4 Method
Figure 2: Overview of the proposed Dual-Domain Meta-Learning framework. The architecture integrates temporal and spectral information by separately encoding input signals through dual-domain multi-prototype learners, which construct diverse representations across domains. These representations are then dynamically modulated by the Statistical Guidance Network (SGN), which derives instance-specific guidance signals and domain importance weights based on statistical characteristics of the input. Finally, a Collaborative Decision module leverages a query-aware attention mechanism to adaptively aggregate class prototypes and refine the classification decision by combining domain-specific logits, inter-prototype relational features, and confidence scores.
@@ 4.2 Dual-Domain Multi-Prototype Meta-Learner
To exploit complementary information, we extract features of fiber data in the time and frequency domains, respectively. Let $x^{(t)}$ and $x^{(f)}$ denote the original time-domain signal and its time-spectrogram of the sample, respectively. We use two parallel feature extractors, $f_{t}(\cdot)$ and $f_{f}(\cdot)$, where the 1DCNN extracts time-domain signal features and the 2DCNN extracts signal time-spectrogram features. They encode $x^{(t)}$ and $x^{(f)}$ into $D$-dimensional feature vectors, $\mathbf{z}^{(t)}=f_{t}(x^{(t)})\in\mathbb{R}^{D}$ and $\mathbf{z}^{(f)}=f_{f}(x^{(f)})\in\mathbb{R}^{D}$. These embedding networks are designed to produce comparable feature representations for both domains (of dimension $D=128$ in our implementation). Given a labeled support set, we obtain a set of support embeddings in the time domain $\mathbf{z}{i}^{(t)}$} and the frequency domain $\mathbf{z}{i}^{(f)}$}. These two domains are handled independently by the same meta-learning logic, as described below.
@@
After obtaining prototypes for each class in both domains, we compute the matching score between each query sample and the class prototype. Specifically, for each query sample $\mathbf{z}_{q}$, we use cosine similarity to measure the similarity between the query and each prototype. For domain $d\in{t,f}$, we calculate the similarity between the query embedding $\mathbf{z}q^{(d)}$ and each prototype $\mathbf{c}{c,k}^{(d)}$ of class $c$. The matching score of class $c$ in domain $d$ is calculated as follows:
@@ 4.3 Statistical Guidance Network (SGN)
The Statistical Guidance Network (SGN) plays a crucial role in our framework, providing context-aware guidance signals to regulate the learning process. It does this by analyzing the global statistical features of the input data. These features, including average energy, zero-crossing rate, and other relevant statistics, are used to guide the meta-learner, provide domain-specific weights, and enhance the prototype matching process.The specific structure is shown in Figure 3.
The SGN receives as input a 26-dimensional statistical feature vector $\mathbf{s}$ for each sample. This vector encodes various global properties of the signal, capturing information not immediately apparent from the original time-domain or frequency-domain representations. The network consists of two sub-networks: one for extracting physical features and the other for extracting environmental context. Each sub-network processes the statistical feature vector $\mathbf{s}$ and outputs latent vectors $\mathbf{p}$ and $\mathbf{e}$, representing the physical and environmental characteristics of the signal, respectively.

### v2

@@ Abstract
Distributed Fiber Optic Sensing (DFOS) is promising for long-range perimeter security, yet practical deployment faces three key obstacles: severe cross-deployment domain shift, scarce or unavailable labels at new sites, and limited within-class coverage even in source deployments. We propose DUPLE, a prototype-based meta-learning framework tailored for cross-deployment DFOS recognition. The core idea is to jointly exploit complementary time- and frequency-domain cues and adapt class representations to sample-specific statistics: (i) a dual-domain learner constructs multi-prototype class representations to cover intra-class heterogeneity; (ii) a lightweight statistical guidance mechanism estimates the reliability of each domain from raw signal statistics; and (iii) a query-adaptive aggregation strategy selects and combines the most relevant prototypes for each query. Extensive experiments on two real-world cross-deployment benchmarks demonstrate consistent improvements over strong deep learning and meta-learning baselines, achieving more accurate and stable recognition under label-scarce target deployments.
@@ 1 Introduction
  • We design a dual-domain meta-learner that jointly encodes time-domain and frequency-domain embeddings and performs per-class multi-prototype clustering in each view. Class evidence from the two domains is fused via a soft-OR (log-sum-exp) rule, yielding expressive and deployment-robust class representations under distribution shifts.
  • We propose SGN to exploit raw statistical descriptors and infer both domain importance and prototype sensitivity. These data-driven priors coordinate the two views and stabilize meta-learning in unlabeled or previously unseen deployments.
  • In the collaborative decision stage, each query uses SGN guidance to form a key and applies attention over the corresponding class-specific prototype sets, adaptively selecting and weighting the most relevant prototypes before producing the final logits. This query-aware aggregation improves cross-deployment generalization with limited supervision.
Extensive experiments on two real-world cross-deployment DFOSdatasets show that the proposed framework consistently outperforms strong deep learning and meta-learning baselines, achieving more accurate and balanced activity recognition across diverse deployment configurations. By aligning the training objective with cross-deployment use, injecting sample-wise statistical priors to guide dual-domain fusion, and enabling instance level adaptation at decision time, these designs deliver stronger generalization with minimal data requirements.
@@ 2 Related Work
To address the domain shift challenge, Domain Generalization (DG) techniques have been proposed to learn domain-invariant representations from multiple source domains. Feature alignment methods, such as Deep CORAL (Sun and Saenko, 2016), align the second-order statistics of deep features between source and target distributions to reduce distributional discrepancies. Alternatively, robust optimization strategies like GroupDRO (Group Distributionally Robust Optimization) (Sagawa et al., 2019) explicitly optimize the model’s performance on the worst-case domain or group, preventing the model from relying on spurious correlations that do not hold across environments. While these DG methods significantly enhance robustness compared to standard empirical risk minimization, they typically aim for a static invariant model and lack the flexibility to actively adapt to specific target deployments using scarce support samples, which is the specific strength of meta-learning frameworks.
@@ 4 Method
Figure 2: Overview of the proposed Dual-Domain Meta-Learning framework. The architecture first integrates temporal and spectral information via a Feature Extraction module, using parallel 1D-CNN and 2D-CNN. Based on these features, the system performs Dual-Domain Prototype Construction to capture intra-class diversity from the support set. In parallel, the SGN Module derives instance-specific guidance signals from the input’s statistical characteristics. This pipeline culminates in the Collaborative Decision module, which leverages the SGN guidance and a query-aware attention mechanism to adaptively aggregate prototypes for the final classification.
@@ 4.2 Dual-Domain Multi-Prototype Meta-Learner
To exploit complementary information, we extract features of fiber data in the time and frequency domains, respectively. Let $x^{(t)}$ and $x^{(f)}$ denote the original time-domain signal and its time-spectrogram of the sample, respectively.
In our implementation, the frequency-domain input $x^{(f)}$ is constructed by Short-Time Fourier Transform (STFT). Given a raw 1-D signal segment sampled at $f_{s}=50\,\mathrm{kHz}$, we compute STFT using a Hann window with FFT size $n_{\mathrm{fft}}=1024$ and hop length $h=512$ (i.e., 50% overlap). Let $Z\in\mathbb{C}^{F\times T}$ denote the complex STFT; we take the magnitude spectrogram and apply logarithmic compression: $x^{(f)}=\log(|Z|+\epsilon)$ with $\epsilon=10^{-8}$. We then apply per-spectrogram standardization to zero mean and unit variance to stabilize the dynamic range. Finally, the spectrogram is padded/cropped to a fixed resolution of $1\times 513\times 294$ (channel $\times$ frequency bins $\times$ time frames) before feeding into the frequency encoder.
We use two parallel feature extractors, $f_{t}(\cdot)$ and $f_{f}(\cdot)$, where the 1DCNN extracts time-domain signal features and the 2DCNN extracts signal time-spectrogram features. They encode $x^{(t)}$ and $x^{(f)}$ into $D$-dimensional feature vectors, $\mathbf{z}^{(t)}=f_{t}(x^{(t)})\in\mathbb{R}^{D}$ and $\mathbf{z}^{(f)}=f_{f}(x^{(f)})\in\mathbb{R}^{D}$. These embedding networks are designed to produce comparable feature representations for both domains (of dimension $D=128$ in our implementation). Given a labeled support set, we obtain a set of support embeddings in the time domain $\mathbf{z}{i}^{(t)}$} and the frequency domain $\mathbf{z}{i}^{(f)}$}. These two domains are handled independently by the same meta-learning logic, as described below.
@@
For each class $c$ in each domain $d\in\{t,f\}$, we determine the number of prototypes $K_{c}^{(d)}$ in a data-driven manner. Let $\mathcal{Z}_{c}^{(d)}=\{\mathbf{z}_{i}^{(d)}\mid y_{i}=c\}$ be the support embeddings of class $c$ and $N_{c}=|\mathcal{Z}_{c}^{(d)}|$. We run $K$-means clustering on $\mathcal{Z}_{c}^{(d)}$ with candidate $K\in\{1,\ldots,\min(K_{\max},N_{c})\}$ and obtain $K$ centroids as prototypes. To select $K_{c}^{(d)}$, we compute the within-cluster sum of squares $\mathrm{WCSS}(K)$ and minimize a penalized criterion:
$J(K)=\mathrm{WCSS}(K)+\alpha\cdot K\cdot\log(\max(2,N_{c})),$ (1)
where $\alpha$ controls the complexity penalty.The selected $K_{c}^{(d)}=\arg\min_{K}J(K)$ balances intra-class multi-modality and over fragmentation.In our experiments, we set $K_{\max}=3$ and $\alpha=0.003$.
After obtaining prototypes for each class in both domains, we compute the matching score between each query sample and the class prototype. Specifically, for each query sample $\mathbf{z}_{q}$, we use cosine similarity to measure the similarity between the query and each prototype. For domain $d\in{t,f}$, we calculate the similarity between the query embedding $\mathbf{z}q^{(d)}$ and each prototype $\mathbf{c}_{c,k}^{(d)}$ of class $c$. The matching score of class $c$ in domain $d$ is calculated as follows:
@@
Table 3: Time-domain statistical features.
Feature Formula Feature Formula
Mean ($S_{1}$) $\mu=\frac{1}{N}\sum_{n=1}^{N}x_{n}$ Std. dev. ($S_{2}$) $\sigma=\sqrt{\frac{1}{N}\sum_{n=1}^{N}(x_{n}-\mu)^{2}}$
Variance ($S_{3}$) $\sigma^{2}=\frac{1}{N}\sum_{n=1}^{N}(x_{n}-\mu)^{2}$ Maximum ($S_{4}$) $\max_{n}x_{n}$
Minimum ($S_{5}$) $\min_{n}x_{n}$ Peak-to-peak ($S_{6}$) $\max_{n}x_{n}-\min_{n}x_{n}$
RMS ($S_{7}$) $\sqrt{\frac{1}{N}\sum_{n=1}^{N}x_{n}^{2}}$ Skewness ($S_{8}$) $\mathrm{skew}(x)=\frac{1}{N}\sum_{n=1}^{N}\big(\frac{x_{n}-\mu}{\sigma}\big)^{3}$
Kurtosis ($S_{9}$) $\mathrm{kurt}(x)=\frac{1}{N}\sum_{n=1}^{N}\big(\frac{x_{n}-\mu}{\sigma}\big)^{4}$ Mean abs. value ($S_{10}$) $\frac{1}{N}\sum_{n=1}^{N}\lvert x_{n}\rvert$
Median ($S_{11}$) $\operatorname{median}(\{x_{n}\})$ IQR ($S_{12}$) $Q_{3}-Q_{1}$
Signal energy ($S_{13}$) $E=\sum_{n=1}^{N}x_{n}^{2}$ Average energy ($S_{14}$) $\frac{E}{N}$
Zero-crossing rate ($S_{15}$) $\mathrm{ZCR}=\frac{N_{\text{zc}}}{N}$ Peak density ($S_{16}$) $\mathrm{PD}=\frac{N_{\text{peaks}}}{N}$
Table 4: Frequency-domain statistical features.
Feature Formula Feature Formula
Spec. mean mag. ($S_{17}$) $\frac{1}{K}\sum_{k=1}^{K}\lvert X(f_{k})\rvert$ Spec. std. ($S_{18}$) $\sqrt{\frac{1}{K}\sum_{k=1}^{K}\big(\lvert X(f_{k})\rvert-\bar{A}\big)^{2}}$
Spec. max mag. ($S_{19}$) $\max_{k}\lvert X(f_{k})\rvert$ Spec. min mag. ($S_{20}$) $\min_{k}\lvert X(f_{k})\rvert$
Spectral centroid ($S_{21}$) $f_{c}=\dfrac{\sum_{k}f_{k}\lvert X(f_{k})\rvert}{\sum_{k}\lvert X(f_{k})\rvert}$ Spectral bandwidth ($S_{22}$) $\sqrt{\dfrac{\sum_{k}(f_{k}-f_{c})^{2}\lvert X(f_{k})\rvert}{\sum_{k}\lvert X(f_{k})\rvert}}$
Spectral roll-off ($S_{23}$) $\sum_{f_{k}\leq f_{r}}\lvert X(f_{k})\rvert^{2}\geq 0.95\sum_{k}\lvert X(f_{k})\rvert^{2}$ Spectral flatness ($S_{24}$) $\dfrac{\exp\big(\frac{1}{K}\sum_{k=1}^{K}\ln\lvert X(f_{k})\rvert\big)}{\frac{1}{K}\sum_{k=1}^{K}\lvert X(f_{k})\rvert}$
Dominant frequency ($S_{25}$) $f_{\text{dom}}=f_{k^{\ast}},\ k^{\ast}=\arg\max_{k}\lvert X(f_{k})\rvert$ Spectral entropy ($S_{26}$) $H=-\sum_{k}p_{k}\log_{2}p_{k},\ p_{k}=\dfrac{\lvert X(f_{k})\rvert}{\sum_{j}\lvert X(f_{j})\rvert}$
@@ 4.3 Statistical Guidance Network (SGN)
The Statistical Guidance Network (SGN) plays a crucial role in our framework, providing context-aware guidance signals to regulate the learning process by analyzing global statistical features of the input data. Concretely, for each signal we extract a 26-dimensional statistical feature vector $\mathbf{s}$ composed of 16 time-domain features and 10 frequency-domain features . The definitions of these time-domain and frequency-domain features are summarized in Tables 3 and 4, respectively. These statistics capture global amplitude, energy and spectral characteristics that are not immediately apparent from the raw time-series or STFT representations, and are used to guide the meta-learner, provide domain-specific weights, and enhance the prototype matching process. The specific structure of the SGN is shown in Figure 3.
The SGN receives as input the 26-dimensional vector $\mathbf{s}$ for each sample. This vector encodes various global properties of the signal, capturing information beyond the original time-domain or frequency-domain representations. The network consists of two sub-networks: one for extracting physical features and the other for extracting environmental context. Each sub-network processes $\mathbf{s}$ and outputs latent vectors $\mathbf{p}$ and $\mathbf{e}$, representing the physical and environmental characteristics of the signal, respectively.
@@ 4.4 Collaborative Decision with Query-Aware Attention
Table 5: Comparison of Traditional, Deep Learning, and Domain Generalization Baselines on OSDG1 and OSDG2 (in %).
Model OSDG1 (%) OSDG2 (%)
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Xgboost(Chen and Guestrin, 2016) $26.58\pm 16.89$ $19.70\pm 10.87$ $33.52\pm 6.60$ $20.76\pm 10.42$ $56.57\pm 47.65$ $68.51\pm 37.90$ $57.93\pm 48.07$ $57.72\pm 45.78$
SVM(Qu et al., 2010) $26.02\pm 13.95$ $21.45\pm 19.11$ $34.34\pm 3.12$ $23.62\pm 15.58$ $53.74\pm 47.00$ $8.96\pm 48.43 55$$5.30\pm 47.73$ $53.25\pm 46.45$
KNN(Jia et al., 2019a) $37.48\pm 13.83$ $42.54\pm 11.15$ $34.92\pm 10.38$ $29.68\pm 9.92$ $60.26\pm 11.89$ $64.50\pm 17.84$ $63.64\pm 9.45$ $56.52\pm 13.83$
1DCNN(Wu et al., 2019a) $34.64\pm 13.53$ $21.56\pm 1.19$ $35.20\pm 6.59$ $22.64\pm 6.77$ $57.38\pm 16.79$ $41.21\pm 24.31$ $51.77\pm 15.90$ $44.09\pm 22.38$
2DCNN(Xu et al., 2018) $42.15\pm 11.10$ $50.94\pm 2.13$ $45.42\pm 6.12$ $38.32\pm 8.07$ $57.31\pm 15.98$ $68.52\pm 19.53$ $58.39\pm 20.19$ $49.44\pm 24.08$
LSTM(Sherstinsky, 2020) $23.19\pm 11.24$ $15.47\pm 1.65$ $30.57\pm 0.52$ $18.33\pm 1.03$ $67.72\pm 8.27$ $57.36\pm 15.68$ $62.69\pm 9.92$ $57.26\pm 12.27$
Transformer(Vaswani et al., 2017) $2.03\pm 3.55 22$$2.37\pm 3.39$ $27.93\pm 0.81$ $21.41\pm 3.95$ $68.69\pm 15.87$ $70.67\pm 14.38$ $67.41\pm 13.47$ $65.92\pm 15.19$
GroupDRO(Sagawa et al., 2019) $57.82\pm 16.10$ $58.87\pm 16.49$ $52.90\pm 3.12$ $48.22\pm 10.46$ $71.72\pm 14.23$ $75.13\pm 22.09$ $74.78\pm 12.36$ $68.15\pm 18.23$
Coral(Sun and Saenko, 2016) $52.21\pm 11.53$ $57.69\pm 6.97$ $9.81\pm 3.30 44$$4.32\pm 13.16$ $70.10\pm 12.25$ $85.65\pm 3.69$ $71.07\pm 6.89$ $64.43\pm 12.67$
Table 6: Few-Shot Accuracy Comparison on OSDG1 and OSDG2 Datasets (in %).
Model OSDG1 (Shots, %) OSDG2 (Shots, %)
1-shot 3-shot 5-shot 10-shot 1-shot 3-shot 5-shot 10-shot
ProtoNet(Snell et al., 2017) $32.71\pm 4.93$ $31.51\pm 5.57$ $31.55\pm 5.84$ $30.80\pm 6.88$ $5.09\pm 11.65 66$$65\pm 11.55 66.$$49\pm 10.77 66.$$6.67\pm 10.22$
MAML(Finn et al., 2017) $38.89\pm 11.39$ $39.76\pm 12.63$ $39.78\pm 12.13$ $39.66\pm 11.87$ $62.56\pm 15.59$ $63.86\pm 13.69$ $64.64\pm 13.73$ $64.87\pm 13.86$
ANIL(Raghu et al., 2019) $30.95\pm 7.54$ $35.15\pm 11.99$ $35.34\pm 11.87$ $36.31\pm 13.42$ $58.87\pm 21.49$ $62.61\pm 18.63$ $63.31\pm 19.08$ $63.67\pm 20.21$
R2D2(Bertinetto et al., 2018) $41.34\pm 8.81$ $40.49\pm 8.09$ $40.24\pm 8.04$ $40.18\pm 7.88$ $54.26\pm 20.19$ $63.34\pm 22.98$ $64.00\pm 23.91$ $64.25\pm 24.74$
FEAT(Ye et al., 2020) $47.29\pm 6.00$ $49.11\pm 9.82$ $49.53\pm 9.07$ $48.80\pm 9.69$ $53.32\pm 5.80$ $57.43\pm 11.49$ $57.74\pm 12.90$ $58.35\pm 14.69$
DUPLE (Ours) $58.73\pm 2.39$ $62.76\pm 2.65$ $63.07\pm 3.56$ $63.66\pm 3.23$ $68.17\pm 15.34$ $7.06\pm 18.98 77$$45\pm 20.35 77.$$7.20\pm 22.73$